# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ Appendix B Proofs
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly increasing and $\Lambda_{n}\nearrow H(\infty)$. Because the convergence is strictly increasing toward $H(\infty)$, we have $\Lambda_{n}<H(\infty)$ for all $n\geq 1$, ensuring the existence of the equilibrium. Thus, for each $n$, $v_{n}(0)$ is well-defined and uniquely determined by $v_{n}(0)=H^{-1}(\Lambda_{n})$. Consequently, the convergence $\Lambda_{n}\to H(\infty)$ implies that $\lim_{n\to\infty}v_{n}(0)=\infty$. By Theorem 3.2, we have $a_{n}(0)=h(v_{n}(0))\lambda_{0}$. Using the asymptotic property of $h$ established in Lemma A.2, namely $\lim_{x\to\infty}h(x)=\frac{1}{r_{0}}$, we obtain
@@
Furthermore, because $\{\Lambda_{n}\}$ is strictly increasing, $\{v_{n}(0)\}$ is strictly increasing. Combined with the monotonicity of $h$ (see Lemma A.1), it follows that $|a_{n}(0)|=|\lambda_{0}|h(v_{n}(0))$ converges monotonically increasing to $\frac{|\lambda_{0}|}{r_{0}}$.
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ is strictly decreasing and $\Lambda_{n}\searrow 0$. Because $H(\infty)>0$, and $\Lambda_{n}\searrow 0$, there exists an integer $N$ such that for all $n>N$, $\Lambda_{n}<H(\infty)$. Thus, the equilibrium exists for all sufficiently large $n$. For such $n$, we have $v_{n}(0)=H^{-1}(\Lambda_{n})$. Because $H(0)=0$ and $H^{-1}$ is continuous, we have
@@
Recall that $h(0)=\frac{1}{\tilde{\mathbb{E}}[\bm{R}]}$. Taking the limit of $a_{n}(0)$, we have
@@
Finally, because $\{\Lambda_{n}\}$ is strictly decreasing, $\{v_{n}(0)\}$ is strictly decreasing. The strict monotonicity of $h$ implies that $|a_{n}(0)|$ converges monotonically decreasing to $\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$.
@@
For notational simplicity, let $h_{n}=h_{\bm{R}_{n}}$, $H_{n}=H_{\bm{R}_{n}}$, and $v_{n}=v_{\bm{R}_{n}}$ ($n\geq 1$). By Lemma A.3, the sequence $\{h_{n}\}_{n\geq 1}$ converges uniformly to $h$ on every compact subset of $0,\infty)$. Because the limit function $h$ is strictly positive, it follows that $\{H_{n}\}_{n\geq 1}$ converges uniformly to $H$ on every compact subset of $0,\infty)$ as well. By Lemma B.1 below, the sequence of the inverse functions $\{H_{n}^{-1}\}_{n\geq 1}$ converge uniformly to $H^{-1}$ on $0,\Lambda(0)]$.
@@
Let $\mathcal{K}\triangleq\left[0,\max_{t\in[0,T]}|v(t)|+1\right]$. Then $v_{n}(t)\in\mathcal{K}$ for all $t\in[0,T]$ and all sufficiently large $n$. Therefore, the uniform convergence of $\{h_{n}\}_{n\geq 1}$ on $\mathcal{K}$ implies $\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v_{n}(t))|\to 0$. Moreover, $h$ is continuous and therefore uniformly continuous on the compact set $\mathcal{K}$, which combined with the fact that $\sup_{t\in[0,T]}|v_{n}(t)-v(t)|\to 0$ implies $\sup_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|\to 0$. Consequently, by (B.1), we have
@@
Finally, recalling that $a_{n}(t)=h_{n}(v_{n}(t))\lambda(t)$ and $a(t)=h(v(t))\lambda(t)$, we get from (B.2) that $\lim_{n\to\infty}a_{n}(t)=a(t)$ for every $t\in[0,T)$. Furthermore, if $\lambda\in L^{\infty}([0,T])$, then the above convergence is uniform on $[0,T)$.
@@
Next, because $H^{-1}$ is continuous on the compact interval $[0,H(M)]$, it is uniformly continuous. Thus, for any $\varepsilon>0$, there exists $\delta>0$ such that for any $x^{\prime},x^{\prime\prime}\in[0,H(M)]$ with $|x^{\prime}-x^{\prime\prime}|<\delta$, we have
@@
By uniform convergence of $H_{n}\to H$ on $[0,M]$, there exists $N\geq N_{0}$ such that for all $n\geq N$ and all $y\in[0,M]$,
@@
Because $x,H(x_{n})\in[0,H(M)]$ and $x_{n}=H^{-1}(H(x_{n}))$, inequality (B.3) yields
@@
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables satisfying the conditions (1) and (2) in Proposition 5.3. Recall the definition in Proposition 3.8 that $\bm{X}_{i}=\bm{R}_{i}-r_{0}$ and $Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]}$, $x\geq 0$. Then, the following assertions hold:
@@
    For each $i\in\{1,2\}$, there exist constant $M_{i}>0$, such that $Q_{i}(x)\leq M_{i}e^{-\frac{1}{2}\delta_{i}x}$ for any $x\geq 0$. In particular, $Q_{i}(x)\to 0$ as $x\to\infty$.
@@
    For each $i\in\{1,2\}$, the integral $\int_{0}^{\infty}(h_{i}^{-2}(s)-r_{0}^{2})ds$ converges to a finite constant $K_{i}$.
@@
    For any $x_{1},x_{2}\geq 0$ and any $\varepsilon>0$, it holds that $\dfrac{Q_{1}(x_{1})}{Q_{2}(x_{2})}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}x_{1}}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x_{2}}}$, where $C$ is a constant depending on $\varepsilon$.
@@
For $i\in\{1,2\}$. By condition (1) in Proposition 5.3, we have $\mathbb{P}\mathbb{(}\bm{X}_{i}=0)=p_{i}$. The Dominated Convergence Theorem then implies $\lim_{x\to\infty}\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]=p_{i}\in(0,1)$. Hence, there exists $m_{i}>0$ such that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]\geq m_{0}$ for any $x\geq 0$. Moreover, by condition (2) in Proposition 5.3, the support of $\bm{X}_{i}$ on $(0,\infty)$ is contained in $\delta_{i},\infty)$. Therefore,
@@
The desired conclusion then follows by taking $M_{i}\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}]}{m_{i}}$.
@@
Combining this with assertion (1) gives the desired bound with $C\triangleq\frac{M_{1}}{\delta_{2}p_{\varepsilon,2}}$. ∎
Now we return to prove Proposition B.3.
By assertion (1) of Lemma B.2 and (A.2), we have $\lim\limits_{x\to\infty}h_{i}(x)=\frac{1}{r_{0}}$. Because $r_{0}>0$, it follows that the integral $\int_{0}^{\infty}h_{i}^{-2}(x)\mathrm{d}x$ diverges, which implies $H_{i}(\infty)=\infty$. Consequently, Theorem 3.2 (1) ensures the existence of a unique equilibrium $\pi_{i}$ for $\bm{R}_{i}$ for any $T>0$. Based on the relation ${\pi}_{i}=(\sigma^{\top})^{-1}a_{i}$, it suffices to compare $a_{i}$. By (2.5), we have $|a_{i}(0,T)|=h_{i}(v_{i}(0,T))|\lambda(0)|$.
First, we establish the asymptotic relationship between $v_{1}(0,T)$ and $v_{2}(0,T)$ as $T\to\infty$. By Theorem 3.2 (1), we have $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$. It then follows that $\lim_{T\to\infty}v_{i}(0,T)=\infty$, because $H_{i}(\infty)=\infty$ and $\lim_{T\to\infty}\Lambda(0,T)=\infty$. Moreover, the relation $v_{i}(0,T)=H_{i}^{-1}(\Lambda(0,T))$ yields
$\Lambda(0,T)=\int_{0}^{v_{i}(0,T)}(r_{0}^{2}+(h_{i}^{-2}(x)-r_{0}^{2}))\mathrm{d}x=r_{0}^{2}v_{i}(0,T)+\int_{0}^{v_{i}(0,T)}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x.$ (B.4)
Define the tail integral $R_{i}(y)\triangleq\int_{y}^{\infty}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x$, which satisfies $\lim_{y\to\infty}R_{i}(y)=0$ by Lemma B.2 (2). Then (B.4) can be rewritten as
$\Lambda(0,T)=r_{0}^{2}v_{i}(0,T)+K_{i}-R_{i}(v_{i}(0,T)).$ (B.5)
@@
$v_{1}(0,T)=v_{2}(0,T)+\dfrac{K_{2}-K_{1}}{r_{0}^{2}}+\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}.$
Let $\Delta K=\dfrac{K_{2}-K_{1}}{r_{0}^{2}}$ and $\xi(T)=\dfrac{R_{2}(v_{2}(0,T))-R_{1}(v_{1}(0,T))}{r_{0}^{2}}$. Because $v_{i}(0,T)\to\infty$ as $T\to\infty$, the tail difference term $\xi(T)$ tends to 0. Consequently, we obtain the asymptotic relation $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$.
Second, we compare $h_{1}(v_{1}(0,T))$ and $h_{2}(v_{2}(0,T))$. Because $h_{i}(s)=(r_{0}+Q_{i}(s))^{-1}$, this reduces to comparing $Q_{1}(v_{1}(0,T))$ and $Q_{2}(v_{2}(0,T))$. Let $\varepsilon>0$ be a constant satisfying $\delta_{1}-\delta_{2}-\varepsilon>0$. Using the bound established in Lemma B.2 (3), we obtain
$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}v_{1}(0,T)}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)v_{2}(0,T)}}.$
Substituting $v_{1}(0,T)=v_{2}(0,T)+\Delta K+\xi(T)$ into the exponent yields
$\dfrac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\leq Ce^{-\frac{1}{2}\delta_{1}(\Delta K+\xi(T))}\cdot e^{-\frac{1}{2}(\delta_{1}-\delta_{2}-\varepsilon)v_{2}(0,T)}.$
Because $\lim_{T\to\infty}\xi(T)=0$, the first exponential term converges to a finite positive constant. Moreover, as $v_{2}(0,T)\to\infty$, the second term tends to 0, and hence $\frac{Q_{1}(v_{1}(0,T))}{Q_{2}(v_{2}(0,T))}\to 0$ as $T\to\infty$.
This implies for sufficiently large $T$, $Q_{1}(v_{1}(0,T))<Q_{2}(v_{2}(0,T))$. Because $h_{i}(x)=(r_{0}+Q_{i}(x))^{-1}$ is strictly decreasing in $Q_{i}$, we have $h_{1}(v_{1}(0,T))>h_{2}(v_{2}(0,T))$. Consequently, $|a_{1}(0,T)|>|a_{2}(0,T)|$ because $\lambda(0)\neq 0$, and equivalently, $|\pi_{1}(0,T)|>|\pi_{2}(0,T)|$.
@@
For $i=1,2$, let $\mathbb{P}(\bm{R}_{i}=r_{0})=p_{i}$. The function $Q_{i}(\cdot)$ is explicitly given by
@@
We first verify the existence of a solution to $|a_{1}(t)|=|a_{2}(t)|$. This equality is equivalent to $Q_{1}(v_{1}(t))=Q_{2}(v_{2}(t))$, because $|a_{i}(t)|=|\lambda(t)|(r_{0}+Q_{i}(v_{i}(t)))^{-1}$. Define $\mathcal{D}(t)\triangleq Q_{1}(v_{1}(t))-Q_{2}(v_{2}(t))$. At $t=0$, by assumption, $\mathcal{D}(0)<0$. At $t=T$, because $v_{i}(T)=0$ and $Q_{i}(0)=\tilde{\mathbb{E}}[\bm{R}_{i}]-r_{0}$, the condition $\tilde{\mathbb{E}}[\bm{R}_{1}]>\tilde{\mathbb{E}}[\bm{R}_{2}]$ implies $Q_{1}(0)>Q_{2}(0)$, and therefore $\mathcal{D}(T)>0$. Because $\mathcal{D}(t)$ is continuous, there exists at least one $t^{*}\in(0,T)$ such that $\mathcal{D}(t^{*})=0$.
Next, we show uniqueness by differentiating $\mathcal{D}(\cdot)$. To this end, we compute $Q_{i}^{\prime}(x)$:

### v3

@@ Appendix B Proofs
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly increases to $H(\infty)$. First, we have that $\Lambda_{n}<H(\infty)$ for all $n\geq 1$, which guarantees the existence of the equilibrium. Then, for each $n$, $v_{n}(0)$ is well-defined and uniquely determined by $v_{n}(0)=H^{-1}(\Lambda_{n})$. Consequently, the convergence $\Lambda_{n}\to H(\infty)$ implies that $\lim_{n\to\infty}v_{n}(0)=\infty$. Based on Theorem 3.2, we have $a_{n}(0)=h(v_{n}(0))\lambda_{0}$. Using the asymptotic property of $h$ established in Lemma A.2, i.e., $\lim_{x\to\infty}h(x)=\frac{1}{r_{0}}$, we obtain
@@
Furthermore, $\{v_{n}(0)\}$ is strictly increasing as $\{\Lambda_{n}\}$ is strictly increasing. Combined with the monotonicity of $h$ (see Lemma A.1), it follows that $|a_{n}(0)|=|\lambda_{0}|h(v_{n}(0))$ increases to $\frac{|\lambda_{0}|}{r_{0}}$.
Assume that the sequence $\{\Lambda_{n}\}_{n\geq 1}$ strictly decreases to 0. By $H(\infty)>0$, $\Lambda_{n}<H(\infty)$ for all sufficiently large $n$. Then the equilibrium exists for all sufficiently large $n$. For such $n$, we have $v_{n}(0)=H^{-1}(\Lambda_{n})$. Because $H(0)=0$ and $H^{-1}$ is continuous, we have
@@
Recall that $h(0)=\frac{1}{\tilde{\mathbb{E}}[\bm{R}]}$. Taking the limit of $a_{n}(0)$ yields
@@
Finally, $\{v_{n}(0)\}$ is strictly decreasing as $\{\Lambda_{n}\}$ is strictly decreasing. The strict monotonicity of $h$ implies that $|a_{n}(0)|$ decreases to $\frac{|\lambda_{0}|}{\tilde{\mathbb{E}}[\bm{R}]}$.
@@
For notational simplicity, let $h_{n}=h_{\bm{R}_{n}}$, $H_{n}=H_{\bm{R}_{n}}$, and $v_{n}=v_{\bm{R}_{n}}$ ($n\geq 1$). By Lemma A.3, the sequence $\{h_{n}\}_{n\geq 1}$ converges uniformly to $h$ on every compact subset of $0,\infty)$. Because the limit function $h$ is strictly positive, it follows that $\{H_{n}\}_{n\geq 1}$ converges uniformly to $H$ on every compact subset of $0,\infty)$ as well. By Lemma B.1 below, the sequence of the inverse functions $\{H_{n}^{-1}\}_{n\geq 1}$ converges uniformly to $H^{-1}$ on $0,\Lambda(0)]$.
@@
Let $\mathcal{K}\triangleq\left[0,\max_{t\in[0,T]}|v(t)|+1\right]$. Then $v_{n}(t)\in\mathcal{K}$ for all $t\in[0,T]$ and all sufficiently large $n$. Therefore, the uniform convergence of $\{h_{n}\}_{n\geq 1}$ on $\mathcal{K}$ implies $\sup_{t\in[0,T]}|h_{n}(v_{n}(t))-h(v_{n}(t))|\to 0$. Moreover, $h$ is continuous and therefore uniformly continuous on the compact set $\mathcal{K}$, which combined with the fact that $\sup\limits_{t\in[0,T]}|v_{n}(t)-v(t)|\to 0$ implies $\sup\limits_{t\in[0,T]}|h(v_{n}(t))-h(v(t))|\to 0$. Consequently, by (B.1), we have
@@
Finally, recalling that $a_{n}(t)=h_{n}(v_{n}(t))\lambda(t)$ and $a(t)=h(v(t))\lambda(t)$, we get from (B.2) that $\lim\limits_{n\to\infty}a_{n}(t)=a(t)$ for every $t\in[0,T)$. Furthermore, if $\lambda\in L^{\infty}([0,T))$, then the above convergence is uniform on $[0,T)$.
@@
Next, because $H^{-1}$ is continuous on the compact interval $[0,H(M)]$, it is uniformly continuous. Thus, for any $\varepsilon>0$, there exists a $\delta>0$ such that for any $x^{\prime},x^{\prime\prime}\in[0,H(M)]$ with $|x^{\prime}-x^{\prime\prime}|<\delta$, we have
@@
By uniform convergence of $H_{n}\to H$ on $[0,M]$, there exists a $N\geq N_{0}$ such that for all $n\geq N$ and all $y\in[0,M]$,
@@
Because $x,H(x_{n})\in[0,H(M)]$ and $x_{n}=H^{-1}(H(x_{n}))$, the inequality (B.3) yields
@@
Let $\bm{R}_{1}$ and $\bm{R}_{2}$ be two non-negative random variables satisfying Conditions (c1) and (c2) in Proposition 5.3. Following the notation in the proof of Lemma A.2, let
$\bm{X}_{i}=\bm{R}_{i}-r_{0}\text{ and }Q_{i}(x)=\frac{\tilde{\mathbb{E}}[\bm{X}_{i}e^{-\frac{1}{2}\bm{X}_{i}x}]}{\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]},\quad x\geq 0.$ (B.4)
Then we have the following assertions.
@@
    For each $i\in\{1,2\}$, there exists a constant $M_{i}>0$ such that $Q_{i}(x)\leq M_{i}e^{-\frac{1}{2}\delta_{i}x}$ for $x\geq 0$. In particular, $Q_{i}(x)\to 0$ as $x\to\infty$.
@@
    For each $i\in\{1,2\}$, the integral $K_{i}=\displaystyle\int_{0}^{\infty}(h_{i}^{-2}(s)-r_{0}^{2})\mathrm{d}s$ is finite.
@@
    For any $\varepsilon>0$, there exists a constant $C>0$ such that $\dfrac{Q_{1}(x_{1})}{Q_{2}(x_{2})}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}x_{1}}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)x_{2}}}$ for any $x_{1},x_{2}\geq 0$.
@@
For $i\in\{1,2\}$. By Condition (c1) in Proposition 5.3, we have $\tilde{\mathbb{P}}\mathbb{(}\bm{X}_{i}=0)=p_{i}$. The dominated convergence theorem then implies $\lim\limits_{x\to\infty}\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]=p_{i}\in(0,1)$. Hence, there exists $m_{i}>0$ such that $\tilde{\mathbb{E}}[e^{-\frac{1}{2}\bm{X}_{i}x}]\geq m_{i}$ for any $x\geq 0$. Moreover, by Condition (c2) in Proposition 5.3, the support of $\bm{X}_{i}$ on $(0,\infty)$ is contained in $\delta_{i},\infty)$. As such,
@@
Thus the desired conclusion follows by taking $M_{i}\triangleq\frac{\tilde{\mathbb{E}}[\bm{X}_{i}\mathbb{I}_{\{\bm{X}_{i}\geq\delta_{i}\}}]}{m_{i}}$.
@@
Combining this with Assertion (1) gives the desired bound with $C\triangleq\frac{M_{1}}{\delta_{2}p_{\varepsilon,2}}$. ∎ Now we return to prove Proposition 5.3.
By Assertion (1) of Lemma B.2 and (A.2), we have $\lim_{x\to\infty}h_{i}(x)=\frac{1}{r_{0}}$. Because $r_{0}>0$, it follows that $\int_{0}^{\infty}h_{i}^{-2}(x)\mathrm{d}x=\infty$, which implies $H_{i}(\infty)=\infty$. Consequently, Theorem 3.2(1) guarantees the existence and uniqueness of the deterministic equilibrium strategy corresponding to $\bm{R}_{i}$ for any deterministic and right-continuous market price of risk $\lambda(\cdot)$ in $L^{2}([0,T))$.
First, we establish the asymptotic relationship between the inverse functions $H_{1}^{-1}(\Lambda)$ and $H_{2}^{-1}(\Lambda)$ for large scalar $\Lambda$. Because $H_{i}(\infty)=\infty$, it follows that $H_{i}^{-1}(\Lambda)\to\infty$ as $\Lambda\to\infty$. Moreover, the identity $\Lambda=H_{i}(H_{i}^{-1}(\Lambda))$ yields
$\Lambda=\int_{0}^{H_{i}^{-1}(\Lambda)}(r_{0}^{2}+(h_{i}^{-2}(x)-r_{0}^{2}))\mathrm{d}x=r_{0}^{2}H_{i}^{-1}(\Lambda)+\int_{0}^{H_{i}^{-1}(\Lambda)}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x.$ (B.5)
Define the tail integral $\Theta_{i}(u)\triangleq\int_{u}^{\infty}(h_{i}^{-2}(x)-r_{0}^{2})\mathrm{d}x$, which satisfies $\lim_{u\to\infty}\Theta_{i}(u)=0$ based on Lemma B.2(2). Then (B.5) can be rewritten as
$\Lambda=r_{0}^{2}H_{i}^{-1}(\Lambda)+K_{i}-\Theta_{i}(H_{i}^{-1}(\Lambda)).$ (B.6)
@@
$H_{1}^{-1}(\Lambda)=H_{2}^{-1}(\Lambda)+\dfrac{K_{2}-K_{1}}{r_{0}^{2}}+\dfrac{\Theta_{2}(H_{2}^{-1}(\Lambda))-\Theta_{1}(H_{1}^{-1}(\Lambda))}{r_{0}^{2}}.$
Let $\Delta K\triangleq\dfrac{K_{2}-K_{1}}{r_{0}^{2}}$ and $\xi(\Lambda)\triangleq\dfrac{\Theta_{2}(H_{2}^{-1}(\Lambda))-\Theta_{1}(H_{1}^{-1}(\Lambda))}{r_{0}^{2}}$. Because $H_{i}^{-1}(\Lambda)\to\infty$ as $\Lambda\to\infty$, the tail difference term $\xi(\Lambda)$ tends to 0 as $\Lambda\to\infty$.
Second, we compare $h_{1}(H_{1}^{-1}(\Lambda))$ and $h_{2}(H_{2}^{-1}(\Lambda))$. Because $h_{i}(s)=(r_{0}+Q_{i}(s))^{-1}$, this reduces to comparing $Q_{1}(H_{1}^{-1}(\Lambda))$ and $Q_{2}(H_{2}^{-1}(\Lambda))$. Let $\varepsilon>0$ be a constant satisfying $\delta_{1}-\delta_{2}-\varepsilon>0$. Using the bound established in Lemma B.2(3), we obtain
$\dfrac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}\leq C\dfrac{e^{-\frac{1}{2}\delta_{1}H_{1}^{-1}(\Lambda)}}{e^{-\frac{1}{2}(\delta_{2}+\varepsilon)H_{2}^{-1}(\Lambda)}}.$
Substituting $H_{1}^{-1}(\Lambda)=H_{2}^{-1}(\Lambda)+\Delta K+\xi(\Lambda)$ into the exponent yields
$\dfrac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}\leq Ce^{-\frac{1}{2}\delta_{1}(\Delta K+\xi(\Lambda))}\cdot e^{-\frac{1}{2}(\delta_{1}-\delta_{2}-\varepsilon)H_{2}^{-1}(\Lambda)}.$
As $\Lambda\to\infty$, we have $\xi(\Lambda)\to 0$, so the first exponential term converges to a finite and positive constant. Moreover, $H_{2}^{-1}(\Lambda)\to\infty$, so the second term tends to 0. Hence, the ratio $\frac{Q_{1}(H_{1}^{-1}(\Lambda))}{Q_{2}(H_{2}^{-1}(\Lambda))}$ tends to 0 as $\Lambda\to\infty$. It follows that there exists a constant $\overline{\Lambda}>0$ such that, for all $\Lambda>\overline{\Lambda}$, we have $Q_{1}(H_{1}^{-1}(\Lambda))<Q_{2}(H_{2}^{-1}(\Lambda))$ and thus
$h_{1}(H_{1}^{-1}(\Lambda))>h_{2}(H_{2}^{-1}(\Lambda))\quad\text{for all }\Lambda>\overline{\Lambda}.$ (B.7)
Next, we compare the risk exposure magnitudes at time 0. For any market price of risk $\lambda(\cdot)$ satisfying $\lambda(0)\neq 0$ and $\Lambda(0)\triangleq\int_{0}^{T}|\lambda(t)|^{2}\,\mathrm{d}t>\overline{\Lambda}$. We have $v_{i}(0)=H_{i}^{-1}(\Lambda(0))$ according to Theorem 3.2(1). Applying (B.7) with $\Lambda=\Lambda(0)$, we obtain $h_{1}(v_{1}(0))>h_{2}(v_{2}(0))$ because $\Lambda(0)>\overline{\Lambda}$. Then, we conclude that $|a_{1}(0)|>|a_{2}(0)|$ because the risk exposure magnitude satisfies $|a_{i}(0)|=h_{i}(v_{i}(0))|\lambda(0)|$ by (2.5).
Finally, because $v_{i}(T)=0$, we have $h_{i}(v_{i}(T))=\dfrac{1}{\tilde{\mathbb{E}}[R_{i}]}$. The assumption $\tilde{\mathbb{E}}[R_{1}]>\tilde{\mathbb{E}}[R_{2}]$ then implies $h_{1}(0)<h_{2}(0)$. Combined with $\lambda(T-)\neq 0$, we obtain $|a_{1}(T-)|<|a_{2}(T-)|$.
@@
Step 1:
Fix a right-continuous and deterministic function $\lambda(\cdot)\in L^{2}([0,T))$. We show that there exists at most one $t^{*}\in(0,T)$ such that $|a_{1}(t^{*})|=|a_{2}(t^{*}|$. Indeed, let $t^{*}\in(0,T)$ be arbitrary and satisfy $|a_{1}(t^{*})|=|a_{2}(t^{*}|$. Let $\mathcal{D}(t)\triangleq Q_{1}(v_{1}(t))-Q_{2}(v_{2}(t))$. Then, $\mathcal{D}(t^{*})=0$. We now investigate the derivative of $\mathcal{D}(\cdot)$ at $t^{*}$. Recalling the expression for $Q_{i}(x)$ in (B.4), we have
@@
Note that